2x+y+a=0 is a focal chord of the parabola y^2=-8x, then the value of a isA)-4B)4C)2D)-2

Grade:12

2 Answers

Vikas TU
6 years ago
If Equation of (y - k)2 = 4p (x - h), where the focus is (h + p, k).
Here by the comparing with the equation given h=0,k=0 and p=-2.
Hence focus = (-2,0).
By substituting the focus into focal chord
2 (-2) + 0 +a = 0
  • a=4
Atul sahay
6 years ago
Comparing given y^2=-8x with standard equation of parabola y^2=-4axSo, 4a=8Hence,a=2Therefore focus is (-2,0)By substituting the focus into focal chordWe get,2×-2+0+k=0Hence,k=4